A sample of 200 college freshmen was asked how many hours per week they spent playing video games. The following frequency distribution presents the results.

Number of Hours Frequency
1.0-

3.9

20
4.0-

6.9

32
7.0-

9.9

47
10.0-

12.9

28
13.0-

15.9

24
16.0-

18.9

18
19.0-

21.9

12
22.0-

24.9

7
25.0-

27.9

4
28.0-

30.9

8

1) What is the class width?

2) What percentage of students play video games for less than 19 hours per week?

3)What percentage of students play video games for 16 or more hours per week?
Answer #1

Solution :

=> From given information ,

=> total n = 200

1)
=> class width :-

class width is the difference between two consecutive lower class limits or two consecutive class boundaries.

=> The first two lower class limits are 1.0 and 4.0.

=> the class width is = 4.0 - 1.0 = 3.0


2)
=> The students play video games for less than 19 hours per week is

=> 20 + 32 + 47 + 28 + 24 + 18 = 169

=> P(less than 19) = 169/200

= 0.845

= 84.5%

=> The percentage of students play video games for less than 19 hours per week is 84.5%

3)   
=> The students play video games for 16 or more hours per week is

=> 18 + 12 + 7 + 4 + 8 = 49

=> P(16 or more) = 49/200

= 0.245

= 24.5%

=> The percentage of students play video games for 16 or more hours per week is 24.5%